# Designing Robots to Help Women*

### v1

@@ III-A2 Health Challenges
Health problems that disproportionately affect women include breast cancer, gynecological/period-related problems (e.g., premenstrual syndrome (PMS)/premenstrual dysmorphic disorder (PMDD), perimenopause/menopause, dysmenorrhea, endometriosis, ovarian and cervical cancer), complications due to pregnancy (including difficult childbirth and abortion, as well as perinatal depression), eating disorders, anxiety, migraines, osteoporosis and autoimmune diseases [28]. [^14] https://online.regiscollege.edu/online-masters-degrees/online-master-science-nursing/womens-health-nurse-practitioner/resources/health-issues-specific-womens-health/
In general, women in the European Union (EU) report poorer health and poorer mental well-being than men, and are more likely to have health limitations over their lifetime [28]. This might be why women are also more inclined than men to attempt suicide [29]. It also often takes women longer to be diagnosed with illnesses, possibly due to gender bias. [^15] https://www.weforum.org/agenda/2024/02/womens-health-gap-healthcare/
@@ III-A3 Other Day-to-Day Challenges
The UN lists eleven largest hurdles for women’s equality by 2030, which includes lack of women in leadership positions, workplace discrimination, and an imbalance in unpaid care work [30]. For example, the report mentions typical rates of 20-30% of managers being female, women earning globally 51 cents per dollar earned by men, and an estimated extra two hours of unpaid care work each day that women will still be expected to spend by 2050.
Physically, women are also on average smaller, shorter and weaker than men (possessing less muscle mass and bone density and more body fat) [31]. This might not merely affect various day-to-day tasks such as reaching high-up items, carrying heavy objects, and opening tight containers; for example, height also affects perceived dominance, which might explain why most Chief Executive Officers (CEOs) of Fortune 500 companies tend to be tall men [32]. Another difference is often seen in clothing, which has repercussions related to health and crime. Some clothes can be used by criminals as an excuse to commit crimes, such as purses that seem easy to snatch, or revealing clothing such as skirts, which can be seen to invite upskirt photography or rape [33]. In Sweden, perpetrators sometimes also come from countries with different cultures, where they might never have seen any of their female relatives (e.g., mother or sister) naked, or women in light clothes; they might also have never been explicitly told by anyone how things are in Sweden.
As well, certain clothes can be impractical, such as high heels, which can lead to foot pain or tripping, or even merely a lack of clothes with (large) pockets, leading to increased use of handbags, which can be stolen [34]. In some countries, women can be hassled or killed if they do not clothe themselves in a certain way (e.g., by Mutawa "religious police" or relatives, if a hijab is deemed to not be sufficiently set); 15 girls were reportedly burned alive in 2002 in an incident when the Mutawa beat girls trying to escape from a burning school without their hijabs [35]. Women are also more targeted by the fashion and cosmetics industries, which are among the biggest international businesses, and encouraged to spend more time on physical appearance and beauty products (e.g., seeking to avoid tanning in Asia) [36]. As well, products are sometimes sold at higher prices targeting women that only differ in packaging, fragrance, or color (e.g., razors similar to those sold to men, but just in pink color); such gender-based price segmentation has been described as a "pink tax", and has also been observed to some extent in Sweden [37].
@@ III-B1 Robots Helping to Fight Crime
In regard to voyeurism, hidden spy cameras could be detected. For example, Sami et al. describe a smartphone app that uses laser/time of flight to detect hidden cameras [38]. Yu et al. used a thermal camera to detect hidden cameras via heat [39]. Here we also build a prototype drone aimed at dissuading voyeurism by detecting hidden cameras. One difference with the concept for our prototype is that both approaches above in their current form still require a human to think of installing and running an app and to move around an apartment searching for devices themselves, which might not be easy in hard-to-reach or high locations. Another difference is that our proposed approach also involves object detection.
@@
In regard to proximate crimes like DV, rape, molestation, and stalking, robots could seek to prevent woman from being targeted and perpetrators from getting away (i.e., preventing the perpetrator’s ability to carry out an attack and ensuring justice is served). Deterrence could be partially achieved by reducing times in which young women must be alone, e.g., via the company of a robot that can record evidence and call for help. Robots could accompany women if there is a perceived risk, such as in a home where DV has occurred, or if she needs to move through a dangerous part of a city, like a parking lot at night. Robots like drones already could seem threatening, taking off, buzzing loudly, flying erratically, and threatening collision, if someone gets too close. This could be like suddenly summoning a loud barking dog to one’s aid, that can otherwise be turned off and carried easily, and could also help to avoid purse-snatching or upskirting. A robot could also indicate a woman’s personal space by projecting a circle of light around her, or shine a light on a stalker following a woman. One related example in the literature exists, of a "spider dress" designed by Anouk Wipprecht that inflates, extending mechanical arms, by monitoring proximity and a woman’s breathing; another variant releases smoke. [^16] https://medium.com/@intel/is-that-a-spider-on-your-dress-or-are-you-happy-to-see-me-da25075314b9 A robot could also prevent criminals from getting too close not only to a woman, but also to objects; e.g., it could prevent someone from getting close to a woman’s drink to insert a sedative. If a woman in a man’s presence suddenly seems incapacitated, this could also indicate that the robot should call for help or separate them. Intent to attack (e.g., hands balled into fists or hidden, getting close, angry/loud language, etc.) could also be detected.
@@
In a more complex scenario, a remote robot might be alerted (e.g., by detecting nearby victims’ screams [40]), or be sent by authorities. In such a case, advanced capabilities might be required to distinguish true attacks from jokes, skits, or play-fighting, and make judgements about "distinction" and "proportionality": A robot might need to identify which person is the victim and which is the perpetrator–a highly challenging task when multiple people are present. As well, decisions about the level of force to use might include some analysis of force differentials, backdrops and crossfire. Additionally, age detection could also be used to detect child abuse or forced marriage, and flying robots or soft robots could be useful to enter closed buildings in which trafficking might be occurring.
@@
During a crime, the robot could target the assailant from a difficult angle (e.g., from behind or from above for a drone) with pepper spray or laser to dazzle the assailant’s eyes, while emitting loud sounds and bright lights to seek help. As well, the robot could try to mark the attacker or their vehicle with paint, and record license plates, or interfere with their movement. The robot could also provide advice during the attack: e.g., to drop to the ground to be harder to move and to avoid being taken to a "second location".
@@
Robots could also help post-hoc, if an attack could not be prevented. For DV, an AI tool could track hospital records to assess risk, if current laws are changed. Federated learning, e.g., downloading updates to AI models for detecting DV without uploading sensitive or restricted data, could be one way to ensure that private data are not misused. After a rape, a robot with a sterile compartment could somehow try to facilitate rape testing immediately, avoiding waiting time at hospitals, also since robots lack the DNA of a human investigator (there might be less risk of contamination). As part of this, robots could also visually analyze victim’s bodies. Fernandes et al. proposed and evaluated a deep learning approach to identify genital lesions that could be evidence of rape, using a dataset of roughly 400 images collected by the Southern Denmark Sexual Assault Referral Centre (78 from non-consensual and 316 from consensual intercourse) [41]. Robots could also help to search for a hostage or corpse, or track fleeing attackers. In the undesired case that a woman has been murdered, robots are being built to do forensics on corpses, like the Virtobot system [42].
@@ III-B2 Robots Helping to Support Health
To detect health problems, a robot could generate three-dimensional (3D) scans of a person using a Lidar device. For example, scans can already be done using Lidar devices in IPhone Pros or IPads. The benefits are:
  • Current methods sometimes haven’t been updated for a long time and could be upgraded. For example, tape measures are used to measure belly girth for pregnancy, which can be inaccurate.
@@
  • People can measure themselves outside of infrequent doctor visits and have continuous control over their own health data (relating to data democratization and data-driven care). For example, in some parts of Sweden, breasts x-rays can be done every five years, and in between, patients are expected to wait.
@@
In more detail, such 3D scans could be used to detect breast cancer, anorexia, and abnormalities:
  • A change in a breast (e.g. due to a lump) could indicate potential cancer.
  • Weight loss in a short time could indicate anorexia.
  • A scan could help someone to understand that it is normal for women to not all be shaped like photo models or porn stars (i.e., there is much variance). For example, this is a goal in the initiative of Visual Sweden called "Visual Vulva". [^17] https://www.visualsweden.se/en/aktuella-projekt/ar-jag-normal/
@@
Where simply scanning is insufficient, robots could intervene in a more complex manner. For example, robots have been designed to carry out clinical breast examinations. [^18] https://www.bristol.ac.uk/news/2023/october/new-robot-could-help-diagnose-breast-cancer-early.html As well, AI systems have been designed for diagnosing ovarian cancer [43], which can be tackled via robotic surgery [44]. Anxiety and depression (which are sometimes exacerbated by loneliness) could be aided by cuddly companion robots such as Paro [45]. To help women with their periods, a robot could apply heat or massage to minimize cramping, and remind about drinking or offer water to minimize bloating. To further help women with eating disorders, a robot could try to prevent binge eating (e.g., hide food or discourage purchases), provide dietary advice, reinforce a positive body image, stop excessive exercise by positive distraction, or cook with a person.
To help women during pregnancy, a robot could perform duties like a midwife, checking baby position (upside down or not) and heart rate, etc. For women experiencing morning sickness, a robot could detect patterns to identify causes of ill feeling (e.g., if certain foods are a problem) and make healthy food. Various robots capable of making food exist, including a prototype we developed that checked how salty or sweet its cooking was to ensure healthiness. [^19] https://www.youtube.com/watch?v=6eqmtkOm-Tk As well, a robot could detect vomiting, and clean when needed. A robot could also detect fainting and either catch the woman or call for help if needed. The robot could also find places for a woman to rest when walking, like benches in shadow, and fan or cool her.
@@ III-B3 Robots Helping with Other Day-to-Day Challenges
Regarding other day-to-day challenges, exoskeletons or other tools could help to level the playing field, by allowing anyone to be large, tall, and strong. For example, this might allow smaller women to become accepted as leaders, or to be accepted as equals, receiving equal pay, in physically intensive jobs such as construction or policing. As well, smaller women could reach high-up items, carry heavy loads, and open tight packages–and maybe even better defend against attackers, helping them to fight back and be too heavy to knock down or abduct. Furthermore, alloparenting robots capable of raising children in a good way could also help [46]; for example, when pay is unequal due to the fear that women will leave work to take care of children, or women are overwhelmed with unpaid care.
@@ III-B4 Sketches
Figure 2: Sketches: (a) Nazar - A helpful flying "eye" could sense "evil" such as hidden cameras in a woman’s room, (b) Hero - A robotic barrier could seek to ensure that people intending harm cannot enter a woman’s personal space, (c) Midwife - A helper robot can take care of pregnant women, (d) Allo - A robotic baby carriage can try to help mothers by taking care of repetitive child raising tasks, (e) Equa-skeleton - Exoskeletons could create equal conditions
@@ IV PROTOTYPE
Theoretical ideas alone sometimes miss practical realities that can be exposed by prototyping. To gain further insight into one of the scenarios, proposed in the first sketch, a proof-of-concept was created, as shown in Fig. 3, in three steps: Preparation. First, a mock-up environment was created containing some objects with hidden cameras and distractor objects intended to make the detection task more difficult. In total, 22 objects were gathered, a similar number to previous papers: Toilet Paper Roll, Book, PET Bottle, Pouch (Snowman), Toy Car, Toy Mammoth, Pill Bottle, Toy Frog, Wine Box, Lamp, Toy Box, Clock, Toy Horse, Gum Bottle, Toy Egg, Glasses, Banana, Sponge, Sunscreen, Toy Buzzer, Medicine Box, Paper Roll. Some standard Logitech/Plexgear webcams were hidden within five objects: Toilet Paper Roll, Toy Car, Pill Bottle, Wine Box, Lamp. Furthermore, a simple OpenCV program showing the video feed was run for each camera to simulate recording. [^20] https://opencv.org/
@@
Robot Motion. Second, we recorded some footage of the drone, a Ryze Tello, flying in front of the objects. We explored various forms of control, e.g., manually controlled the drone via the Ryze Tello smartphone app, as well as using the EasyTello python library [^21] https://pypi.org/project/easytello/ to issue commands and obtain video, and detecting ArUco markers that could be used to guide the robot or for distance estimation.
@@
Hidden camera detection. Third, we explored how spy cameras could be detected using a thermal camera/RGB camera. Since the cameras weighed too much (90g without battery, vs. 80g payload for the drone), for our initial exploration, data were obtained with the camera placed on a desk approximately 1.5m in front of the mock-up environment.
@@
Fig. 4 shows this basic process: Raw RGB and thermal data were obtained using, respectively, a Sony IMX219 8-megapixel sensor, and an inexpensive 80 x 60 FLIR camera capable of detecting heat in the range of 8–14 µm. After calculating a mapping, the thermal camera was used to find warm areas in view that might arise from a hidden camera. Simplified thresholding was conducted to derive a mask. Next, an algorithm was run to detect which objects might be responsible for the warm spots. The aim was to reduce the area that needs to be searched and make it easier to find the cameras. YOLO version 8s from Ultralytics was run on the RGB image with the confidence parameter set to 0.1 to detect the locations of objects as a set of bounding boxes. [^22] https://www.ultralytics.com/yolo YOLO uses deep learning (a convolutional neural network) to detect objects, where the confidence threshold handles non-maximum suppression. Finally, bounding boxes enclosing heat traces were selected as the output of the system, representing objects potentially concealing cameras, that a robot could either remove or show to a human.
@@
As a result, 5/5 of the cameras’ heat signatures were clearly visible after thermal thresholding, resulting in five detected contours. Next, boundary boxes were detected for 4/5 (80%) of the objects enclosing hidden cameras (although overall, only half, or 11, or the 22 objects were detected by YOLO). Finally, the agreement between the ground truth and system output regarding locations of objects concealing hidden cameras was calculated. The average Intersection over Union (IoU) was 0.401 (min=0.0 for the Toy Car which was not detected as an object, max=0.881 for the Wine Box).
@@
Figure 4: Image Processing: (a-b) raw RGB and thermal images, (c-d) intermediate output from YOLO and threshold on thermal traces, (e-f) comparison of overall groundtruth–where purple boxes with red asterisks are the target, and distractor objects are labelled in other colors–and the system output, with groundtruth in blue, output in green

### v2

@@ III-A2 Health Challenges
Some health problems typically only affect women, such as breast cancer, gynecological problems (e.g., premenstrual syndrome (PMS)/premenstrual dysmorphic disorder (PMDD), perimenopause/menopause, dysmenorrhea, endometriosis, ovarian and cervical cancer), and complications due to pregnancy (including difficult childbirth and abortion, as well as perinatal depression) [29].  [^16] https://online.regiscollege.edu/online-masters-degrees/online-master-science-nursing/womens-health-nurse-practitioner/resources/health-issues-specific-womens-health Other health problems also affect men but disproportionately affect women, such as eating disorders, anxiety, migraines, osteoporosis and autoimmune diseases.
As such, women in the European Union also report poorer health and mental well-being than men, and are more likely to have health limitations over their lifetime [29]. Obtaining a diagnosis also often takes longer for women, possibly due to gender bias.171717weforum.org/agenda/2024/02/womens-health-gap-healthcare This might also play some role in why women are more inclined than men to attempt suicide [30].
@@ III-A3 Other Day-to-Day Challenges
Women also deal with inequalities related to the workplace, physical characteristics, and social expectations regarding appearance: The UN lists eleven largest hurdles for women’s equality by 2030, which includes lack of women in leadership positions, workplace discrimination, and an imbalance in unpaid care work [31]. For example, the report mentions typical rates of 20-30% of managers being female, women earning globally 51 cents per dollar earned by men, and an estimated extra two hours of unpaid care work each day that women will still be expected to spend by 2050.
Physically, women are also on average smaller, shorter and weaker than men (possessing less muscle mass and bone density and more body fat) [32]. This could affect more than day-to-day tasks such as reaching high-up items, carrying heavy objects, and opening tight containers; for example, height also affects perceived dominance, which might explain why most Chief Executive Officers (CEOs) of Fortune 500 companies tend to be tall men [33].
Also regarding appearances, expectations of how women should dress themselves can have repercussions related to crime and health. Fashion items have been mentioned by criminals seeking to justify their crimes: Lacking sufficient pocket space can lead to the use of handbags, which can be seen as easy to snatch and invite theft [34]. Revealing clothing can also be seen to invite upskirt photography or rape [35]. In Sweden, perpetrators sometimes also come from countries with different cultures, where they might never have seen any of their female relatives (e.g., mother or sister) naked or in light clothes, and might never have been heard about norms for dressing in Sweden. In some countries, women can also be hassled or killed if they do not clothe themselves in a certain way (e.g., by Mutawa "religious police" or relatives, if a hijab is deemed to not be sufficiently set); 15 girls were reportedly burned alive in 2002 in an incident when the Mutawa beat girls trying to escape from a burning school without their hijabs [36]. Also, while tight corsets and footbinding are no longer common, high heels can still lead to reduced control when driving, foot pain, or tripping.
Like with clothing, women are also more targeted by the fashion and cosmetics industries, which are among the biggest international businesses, and encouraged to spend more time on physical appearance (e.g., seeking to avoid tanning in Asia) [37]. As well, products are sometimes sold at higher prices targeting women that only differ in packaging, fragrance, or color (e.g., razors similar to those sold to men, but just in pink color); such gender-based price segmentation has been described as a "pink tax", and has also been observed to some extent in Sweden [38].
@@ III-B1 Robots Helping to Fight Crime
Robots could interfere with crimes, by increasing women’s awareness of criminals and making it more difficult for criminals to escape prosecution (i.e., preventing a perpetrator’s ability to carry out an attack and ensuring justice is served). In regard to voyeurism, hidden spy cameras could be detected, e.g., by checking for unusual objects overlooking sensitive areas (e.g., with reflective lights from lenses, wires, or lights), unknown Wi-Fi devices or radio frequency communications, interference with phone signals, or buzzing sounds. [^18] https://reolink.com/blog/how-to-detect-hidden-cameras For example, Sami et al. described a smartphone app that uses lasers/time of flight to detect hidden cameras [39]. Yu et al. conversely used a thermal camera to detect hidden cameras via heat [40]. In the current paper, we also describe a prototype drone aimed at dissuading voyeurism by detecting hidden cameras. Two differences are that the above approaches require a human to run an app and move around an apartment searching for devices themselves, which might not be easy in hard-to-reach or high locations, and that we also consider detecting object boundaries to reduce the search space.
@@
Likewise, in regard to DV and rape (and possibly also to other crimes such as molestation, stalking, upskirting, and purse-snatching), robots could seek to prevent a woman from being targeted and perpetrators from getting away. Deterrence could be partially achieved by reducing times in which women at risk must be alone. For example, a robot capable of recording evidence and calling for help could accompany a woman in risky environments such as a home where DV has occurred, or if she needs to move through a dangerous part of a city, like a parking lot at night. One option could be to deploy a drone from a nearby rooftop, as police in Southern Sweden are exploring. [^19] https://www.svt.se/nyheter/lokalt/skane/p-platser-for-polisens-dronare-byggs-pa-hustak-i-malmo (Swedish) If danger seems likely, a drone could threaten criminals by, e.g., buzzing loudly, flying erratically, moving at head height, and potentially causing injury if a criminal gets too close and causes a collision. This could be like suddenly being able to summon a loud barking dog to one’s aid, with the benefit that a robot could be sent back to its station, or turned off and carried, when the danger is gone. Additionally, a robot could also indicate a woman’s personal space by projecting a circle of light around her, or shine light on a stalker. One related example in the literature exists, of a "spider dress" designed by Anouk Wipprecht that inflates, extending mechanical arms, based on monitoring proximity and a woman’s breathing; another variant releases smoke. [^20] https://medium.com/@intel/is-that-a-spider-on-your-dress-or-are-you-happy-to-see-me-da25075314b9
@@
Furthermore, a more complex robot could seek to also protect certain objects, like a woman’s drink at a bar; detect if a woman suddenly seems incapacitated and in danger of being abducted; or infer an intent to attack (e.g., hands balled into fists or hidden, getting close, angry/loud language, etc.) A remote robot might also be alerted (e.g., by detecting nearby victims’ screams [41]), or be sent by authorities. In such a case, advanced capabilities might be required to distinguish true attacks from jokes, skits, or play-fighting, and make judgements about "distinction" and "proportionality": A robot might need to identify which person is the victim and which is the perpetrator–a highly challenging task when multiple people are present. As well, decisions about the level of force to use might include some analysis of force differentials, backdrops and crossfire. Additionally, age detection could also be used to detect child abuse or forced marriage, and flying robots or soft robots could be useful to enter closed buildings in which trafficking might be occurring.
@@
During a crime, a robot could target the assailant from a difficult angle (e.g., from behind or from above for a drone) with pepper spray or laser to dazzle the assailant’s eyes, while emitting loud sounds and bright lights to seek help. As well, the robot could try to mark the attacker or their vehicle with paint, and record license plates, or interfere with their movement. The robot could also provide advice during an attack: e.g., to drop to the ground to be harder to move and to avoid being taken to a "second location".
@@
Robots could also help post-hoc, if an attack could not be prevented. For DV, an AI tool could track hospital records to assess risk, if current laws change. Federated learning, e.g., downloading updates to detection models without uploading sensitive or restricted data, could also be one way to ensure that private data are not misused. After a rape, a robot with a sterile compartment could try to facilitate rape testing immediately, avoiding waiting time at hospitals–possibly with less risk of contamination, since robots lack the DNA of a human investigator. As part of this, robots could also visually analyze victims’ bodies. For example, Fernandes et al. reported on a deep learning approach that can identify genital lesions indicative of rape, using a dataset of roughly 400 images collected by the Southern Denmark Sexual Assault Referral Centre (78 from non-consensual and 316 from consensual intercourse) [42]. Robots could also help to search for a hostage or corpse, or track fleeing attackers. And, in the undesired case that a woman has been murdered, robots are also being built to do forensics on corpses, like the Virtobot system [43].
@@ III-B2 Robots Helping to Support Health
Robots could also help to detect health problems and intervene, in line with the ideas of democratization of health care and data-driven care. To detect problems, a robot could generate three-dimensional (3D) scans of a person, e.g., using a Lidar or ultrasound device, along with other sensing. Scans already possible today using Lidar devices in iPhone Pros or iPads could be facilitated by robots, that can continuously and accurately scan at arbitrary distances, without requiring human time or effort. Thus, better health outcomes could result by enabling improved methods and continuous measurement:
  • Outdated, suboptimal methods could be replaced, such as using tape measures to measure belly girth for pregnancy, which can be inaccurate; thus, a robot could perform duties like a midwife, checking a baby’s position (upside down or not) in addition to heart rate, etc. As well, AI systems can aid, e.g., in the ultrasound diagnosis of ovarian cancer [44].
@@
  • People could measure themselves outside of infrequent doctor visits and have continuous control over their own health data. For example, in some parts of Sweden, patients wait five years between breast X-rays.
@@
In more detail, 3D scans could be used to detect breast cancer, anorexia, and abnormalities:
In addition to visual scans, robots have been designed to haptically carry out clinical breast examinations. [^23] https://www.bristol.ac.uk/news/2023/october/new-robot-could-help-diagnose-breast-cancer-early.html As well, for women experiencing morning sickness, a robot could seek to detect causes of ill feeling (e.g., if certain foods are a problem), vomiting, fainting, or overheating.
@@
Where simply scanning is insufficient, robots could intervene in a more complex manner, providing healthcare, positive social interactions and touch, information, or sustenance: For example, problems such as ovarian cancer can be tackled with robotic surgery  [45] Anxiety and depression, which are sometimes exacerbated by loneliness, could be aided by positive social interactions with cuddly robots such as Paro [46]. To help women with their periods, a robot could apply heat or massage to minimize cramping, and remind about drinking or offer water to minimize bloating. To help women with eating disorders, a robot could try to prevent binge eating (e.g., hide food or discourage purchases), provide dietary advice, reinforce a positive body image, stop excessive exercise by positive distraction, or cook with a person. Cooking could also help women experiencing morning sickness. (Various robots capable of making food exist, including a prototype we developed that checked how salty or sweet its cooking was to ensure healthiness. [^24] https://www.youtube.com/watch?v=6eqmtkOm-Tk) As well, a robot could clean after vomiting; catch a fainting woman or call for help; and find places for a woman to rest when walking, like benches in shadow, and fan or cool her. A childlike embodiment for such a robot could also suggest that a woman is with child even when her bump is not clearly visible.
@@ III-B3 Robots Helping with Other Day-to-Day Challenges
Robots could also support equality by enhancing capabilities and freeing up time: Exoskeletons or other tools could help to level the playing field, by allowing anyone to be large, tall, and strong. For example, this might allow smaller women to become accepted as leaders, or peers receiving equal pay, in physically intensive jobs such as construction or policing. As well, smaller women could reach high-up items, carry heavy loads, and open tight packages–and maybe even better defend against attackers, helping them to fight back and be too heavy to knock down or abduct. Furthermore, alloparenting robots capable of raising children in a good way could also help [47]–for example, when pay is unequal due to the fear that women will leave work to take care of children, or women are overwhelmed with unpaid care. For example, a wheeled robot could carry children, while playing with them and providing positive attention.
@@ III-B4 Sketches
Figure 2: Sketches: (a) Nazar - A helpful flying "eye" could sense "evil" such as hidden cameras located in high-up or hard-to-see places for women seeking privacy, (b) Hero - A robotic barrier could seek to ensure that people intending harm cannot enter a woman’s personal space in dangerous places, (c) Midwife - A helper robot could seek to maintain good living conditions for pregnant women, (d) Allo - A robotic baby carriage could help mothers by taking care of repetitive child raising tasks, (e) Equa-skeleton - Exoskeletons could create equal conditions for working women
@@ IV PROTOTYPE
Theoretical ideas alone sometimes miss practical realities that can be exposed by prototyping. To gain further insight into one of the scenarios, proposed in the first sketch, a proof-of-concept was created, as shown in Fig. 3, in three steps: Preparation. First, a mock-up environment was created containing some objects with hidden cameras and distractor objects intended to make the detection task more difficult. In total, 22 objects were gathered, a similar number to previous papers: Toilet Paper Roll, Book, PET Bottle, Pouch (Snowman), Toy Car, Toy Mammoth, Pill Bottle, Toy Frog, Wine Box, Lamp, Toy Box, Clock, Toy Horse, Gum Bottle, Toy Egg, Glasses, Banana, Sponge, Sunscreen, Toy Buzzer, Medicine Box, Paper Roll. Some standard Logitech/Plexgear webcams were hidden within five objects: Toilet Paper Roll, Toy Car, Pill Bottle, Wine Box, Lamp. Furthermore, a simple OpenCV program showing the video feed was run for each camera to simulate recording. [^25] https://opencv.org
@@
Robot Motion. Second, we recorded some footage of the drone, a Ryze Tello, flying in front of the objects. We explored various forms of control, e.g., manually controlled the drone via the Ryze Tello smartphone app, as well as using the EasyTello python library [^26] https://pypi.org/project/easytello to issue commands and obtain video, and detecting ArUco markers that could be used to guide the robot or for distance estimation.
@@
Hidden camera detection. Third, we explored how spy cameras could be detected using an RGB/thermal camera. Since the camera weighed too much (90g without battery vs. 80g payload for the drone), data for our initial exploration were obtained with the camera placed on a desk approximately 1.5m in front of the mock-up environment.
@@
Fig. 4 shows this basic process: Raw RGB and thermal data were obtained using, respectively, a Sony IMX219 8-megapixel sensor, and an inexpensive 80 x 60 FLIR camera capable of detecting heat in the range of 8–14 µm. After calculating a mapping, the thermal camera was used to find warm areas in view that might arise from a hidden camera. Simplified thresholding was conducted to derive a mask. Next, an algorithm detected which objects might be responsible for the warm spots, to reduce the area that needs to be searched and make it easier to find the cameras. YOLO version 8s from Ultralytics was run on the RGB image with the confidence parameter set to 0.1 to detect the locations of objects as a set of bounding boxes. [^27] https://www.ultralytics.com/yolo YOLO uses deep learning (a convolutional neural network) to detect objects, where the confidence threshold handles non-maximum suppression. Finally, bounding boxes enclosing heat traces were selected as the output of the system, representing objects potentially concealing cameras, that a robot could either remove or show to a human.
@@
As a result, 5/5 of the cameras’ heat signatures were clearly visible after thermal thresholding, resulting in five detected contours. Boundary boxes were detected for 4/5 (80%) of the objects enclosing hidden cameras (although overall, only half of the 22 objects were detected by YOLO). Finally, the agreement between the ground truth and system output regarding locations of objects concealing hidden cameras was calculated. The average Intersection over Union (IoU) was 0.401 (min=0.0 for the Toy Car which was not detected as an object, max=0.881 for the Wine Box).
@@
Figure 4: Image Processing: (a-b) raw RGB and thermal images, (c-d) intermediate output from YOLO and threshold on thermal traces, (e-f) comparison of overall ground truth on the left–where purple boxes with red asterisks are the target, and distractor objects are labelled in other colors–with the system output on the right in green (ground truth for target objects repeated in blue)